A driver travels northbound on a highway at a spee
Answer #1

Let the speed of the sound in air is 343m/sec.
a) Relative speed of the vehicles = (26 + 32) = 58m/sec.
(343 + 58) = apparent sound speed of 401m/sec. (As they are approcing together)
2600 x (401/343) = 3040Hz. heard.

b) (343 - 58) = apparent sound speed of 285m/sec. (As they are moving away)
2600 x (285/343) = 2160Hz. heard.

c) Relative speeds of the vehicles becomes (32 - 26) = 6m/sec.

While police overtakes = 2600 * (343 + 6) / 343 = 2645 Hz

after police car passes = 2600 * (343 - 6) / 343 = 2554 Hz